# Vision LLMs Are Bad at Hierarchical Visual Understanding, and LLMs Are the Bottleneck

### v1

@@ C.1.3 Taxonomy as Context
C.1.4 Questions with Binary Answer
@@
Statement 1: <image> The bird in the image belongs to the <hierarchy>(e.g., Order) of <ground truth> (e.g., Passeriformes).Is this statement correct? Please answer Yes or No.Statement 2: <image> The bird in the image belongs to the <hierarchy>(e.g., Order) of <similar class>.Is this statement correct? Please answer Yes or No.Statement 3: <image> The bird in the image belongs to the <hierarchy>(e.g., Order) of <similar class>.Is this statement correct? Please answer Yes or No.Statement 4: <image> The bird in the image belongs to the <hierarchy>(e.g., Order) of <similar class>.Is this statement correct? Please answer Yes or $\begin{array}[]{l}\text{{Statement 1:} {<image> The bird in the image belongs to the <hierarchy>}}\\ \text{{(e.g., Order) of <ground truth> (e.g., Passeriformes).}}\\ \text{{Is this statement correct? Please answer Yes or No.}}\\ \text{{Statement 2:} {<image> The bird in the image belongs to the <hierarchy>}}\\ \text{{(e.g., Order) of <similar class>.}}\\ \text{{Is this statement correct? Please answer Yes or No.}}\\ \text{{Statement 3:} {<image> The bird in the image belongs to the <hierarchy>}}\\ \text{{(e.g., Order) of <similar class>.}}\\ \text{{Is this statement correct? Please answer Yes or No.}}\\ \text{{Statement 4:} {<image> The bird in the image belongs to the <hierarchy>}}\\ \text{{(e.g., Order) of <similar class>.}}\\ \text{{Is this statement correct? Please answer Yes or No.}}\end{array}$
@@ C.2 Linear Probing of Visual Features
[…]
Qwen2.5-VL-32B [4] 67.88 72.88 41.91 82.18 90.62 0.9517
Qwen2.5-VL-72B [4] 83.08 87.76 41.19 84.51 92.98 0.9192
@@ C.4 HCA over Different Taxonomy Depth
Qwen2.5-VL-7B [4] 17.67 35.15 51.86 65.32 90.53 98.81
InternVL2.5-8B [9] 5.66 14.58 28.35 43.03 74.82 90.99
LLaVA-OV-7B [29] 4.62 11.83 25.22 42.30 78.40 96.12
@@
Qwen2.5-VL-7B [4] 64.22 60.06 79.78 65.99 99.86 N/A
InternVL2.5-8B [9] 41.15 38.38 65.49 83.76 99.67 N/A
LLaVA-OV-7B [29] 28.49 27.95 55.20 49.46 99.82 N/A
@@ C.5 Comparison Between Vision-Tuned LLMs and Original LLMs
Figure 12: HCA difference between vision-tuned LLMs and their original versions across all 7B/8B open-source VLLMs. ($\Delta$HCA = Vision-Tuned HCA - Original HCA.)
@@ D.4 Evaluation on General VQA Benchmarks
Appendix E Limitations
While we have identified that the bottleneck in VLLM’s hierarchical visual understanding lies in the LLM component, the underlying cause of LLMs’ lack of hierarchical consistency in the language space remains an open question. Given the vast, highly structured corpora used during pre-training, one might expect stronger hierarchical representations to emerge from LLMs naturally. Unfortunately, our computational budget precludes training an LLM from scratch to verify this hypothesis. We, therefore, leave to future work the investigation of pre-training strategies that inject explicit hierarchical knowledge, an avenue that could clarify the underlying cause and potentially close the remaining performance gap.
Moreover, our study focused on hierarchical image classification due to limited resources. However, hierarchical visual understanding is broader, including video, 3D, and other visual modalities and more diverse taxonomies. We conjecture that state-of-the-art VLLMs would still perform poorly in those scenarios, but the causes could be different from our findings. LLMs probably would remain the weak point in those scenarios, and it is possible that the visual encoder or projector would be equally responsible.
Finally, we made a bold hypothesis that one cannot make VLLMs understand visual concepts fully hierarchical until LLMs possess corresponding taxonomy knowledge. It could overly blame LLMs, although we have supported this hypothesis with a systematic empirical investigation and the strong correlations between LLMs’ taxonomy knowledge and the corresponding VLLMs’ hierarchical visual understanding performance. Some post-training and test-time computation methods could work well without explicitly improving LLMs’ taxonomy knowledge.
Appendix F Broader Impacts and Ethics Statement
By pinpointing the LLM component as the bottleneck in VLLM’s hierarchical visual understanding and demonstrating that modest multimodal finetuning already improves textual taxonomy knowledge, our findings encourage the community to (i) incorporate explicit hierarchical objectives during LLM pre-training, (ii) curate multimodal corpora with reliable taxonomic annotations, and (iii) develop evaluation metrics that penalize hierarchical inconsistency. These steps could yield models that are safer and more trustworthy in real-world, hierarchy-rich settings.
Potential downsides include the amplification of existing taxonomic biases or misclassifications if the training data encode culturally or scientifically outdated hierarchies. Researchers should therefore audit hierarchies for regional or disciplinary bias, publish data-curation protocols, and, where feasible, provide mechanisms for community feedback and correction.
Overall, we believe that exposing and remedying hierarchical blind spots in VLLMs will enable more reliable AI systems and support scientific, environmental, and industrial domains that depend on structured semantic understanding.
Appendix G Detailed Discussion of Related Works
@@
Hierarchical classification. Hierarchical classification [47, 25] involves assigning labels from a structured semantic hierarchy rather than from a flat label space lacking relational structure. In the vision domain, hierarchical image classification aims to improve visual consistency across coarse-to-fine categories, thereby enhancing overall classification performance. Recent work has introduced structural priors into visual models through hierarchical loss functions, multi-level supervision, and taxonomy-aligned embeddings [61, 43, 65, 48, 7]. Beyond the visual domain, hierarchical classification has also been extensively explored in the language domain [70, 56, 71]. Similar to approaches developed for enhancing hierarchical consistency in vision models, prior work has focused on injecting hierarchical information into language encoders to improve the structure-awareness of text embeddings. Another line of research aims to understand how hierarchical structures are inherently encoded within pre-trained language models. He et al. [21] retrained transformer-based language models in hyperbolic space, resulting in improved modeling of hierarchical knowledge.

### v2

@@ C.1.3 Taxonomy as Context
C.1.4 Few-shot VQA
We conducted additional experiments on the CUB-200 dataset using Qwen2.5-VL-7B with few-shot prompting ranging from 1 to 5 shots. We use level-specific QA pairs as few-shot examples to evaluate performance at each hierarchy level. An example is provided as follows:
[…]
Results are reported in Table 17. Interestingly, we observe no performance improvement across different numbers of few-shot examples.
Table 17: HCA performance across different few-shot settings.
#Few-shot 0 1 2 3 4 5
HCA 66.26 64.83 65.57 65.50 66.10 65.46
C.1.5 Questions with Binary Answer
@@
Statement 1: <image> The bird in the image belongs to the <hierarchy>
(e.g., Order) of <ground truth> (e.g., Passeriformes).
Is this statement correct? Please answer Yes or No.
Statement 2: <image> The bird in the image belongs to the <hierarchy>
(e.g., Order) of <similar class>.
Is this statement correct? Please answer Yes or No.
Statement 3: <image> The bird in the image belongs to the <hierarchy>
(e.g., Order) of <similar class>.
Is this statement correct? Please answer Yes or No.
Statement 4: <image> The bird in the image belongs to the <hierarchy>
(e.g., Order) of <similar class>.
Is this statement correct? Please answer Yes or No.
@@ C.2 Linear Probing of Visual Features
LLM of iNat21-Animal iNat21-Plant ImgNet-Artifact ImgNet-Animal CUB-200 $\rho(\text{text,visual})$
Qwen2.5-VL-32B 67.88 72.88 41.91 82.18 90.62 0.9517
Qwen2.5-VL-72B 83.08 87.76 41.19 84.51 92.98 0.9192
@@ C.4 HCA over Different Taxonomy Depth
Qwen2.5-VL-7B 17.67 35.15 51.86 65.32 90.53 98.81
InternVL2.5-8B 5.66 14.58 28.35 43.03 74.82 90.99
LLaVA-OV-7B 4.62 11.83 25.22 42.30 78.40 96.12
@@
Qwen2.5-VL-7B 64.22 60.06 79.78 65.99 99.86 N/A
InternVL2.5-8B 41.15 38.38 65.49 83.76 99.67 N/A
LLaVA-OV-7B 28.49 27.95 55.20 49.46 99.82 N/A
@@ C.5 Comparison Between Vision-Tuned LLMs and Original LLMs
[…]
@@ D.4 Evaluation on General VQA Benchmarks
Appendix E Detailed Discussion of Related Works
@@
Hierarchical classification. Hierarchical classification [56, 29] involves assigning labels from a structured semantic hierarchy rather than from a flat label space lacking relational structure. In the vision domain, hierarchical image classification aims to improve visual consistency across coarse-to-fine categories, thereby enhancing overall classification performance. Recent work has introduced structural priors into visual models through hierarchical loss functions, multi-level supervision, and taxonomy-aligned embeddings [73, 53, 77, 57, 9]. Beyond the visual domain, hierarchical classification has also been extensively explored in the language domain [81, 66, 82]. Similar to approaches developed for enhancing hierarchical consistency in vision models, prior work has focused on injecting hierarchical information into language encoders to improve the structure-awareness of text embeddings. He et al. [26] retrained transformer-based language models in hyperbolic space, resulting in improved modeling of hierarchical knowledge. Another line of research aims to understand how hierarchical structures are inherently encoded within pre-trained language models. Nikishina et al. [45] provide a comprehensive analysis of transformer-based models for the task of hypernymy prediction, evaluating their ability to infer IS-A relations. Lin and Ng [37] study whether pre-trained BERT models capture the transitivity of IS-A relations in WordNet. To more rigorously assess such capabilities, He et al. [25] propose ONTOLAMA, a benchmark and evaluation framework targeting subsumption inference within ontologies. Similarly, Moskvoretskii et al. [44] evaluate the WordNet-based lexical-semantic reasoning ability of the LLaMA-2-7B model through the TaxoLLaMA framework.
@@
Appendix F Limitations
While we have identified that the bottleneck in VLLM’s hierarchical visual recognition lies in the LLM component, the underlying cause of LLMs’ lack of hierarchical consistency in the language space remains an open question. Given the vast, highly structured corpora used during pre-training, one might expect stronger hierarchical representations to emerge from LLMs naturally. Unfortunately, our computational budget precludes training an LLM from scratch to verify this hypothesis. We, therefore, leave to future work the investigation of pre-training strategies that inject explicit hierarchical knowledge, an avenue that could clarify the underlying cause and potentially close the remaining performance gap.
Moreover, our study focused on hierarchical image classification due to limited resources. However, hierarchical visual recognition is broader, including video, 3D, and other visual modalities and more diverse taxonomies. We conjecture that state-of-the-art VLLMs would still perform poorly in those scenarios, but the causes could be different from our findings. LLMs probably would remain the weak point in those scenarios, and it is possible that the visual encoder or projector would be equally responsible.
Finally, we made a bold hypothesis that one cannot make VLLMs understand visual concepts fully hierarchical until LLMs possess corresponding taxonomy knowledge. It could overly blame LLMs, although we have supported this hypothesis with a systematic empirical investigation and the strong correlations between LLMs’ taxonomy knowledge and the corresponding VLLMs’ hierarchical visual recognition performance. Some post-training and test-time computation methods could work well without explicitly improving LLMs’ taxonomy knowledge.
Appendix G Broader Impacts
By pinpointing the LLM component as the bottleneck in VLLM’s hierarchical visual recognition and demonstrating that modest multimodal finetuning already improves textual taxonomy knowledge, our findings encourage the community to (i) incorporate explicit hierarchical objectives during LLM pre-training, (ii) curate multimodal corpora with reliable taxonomic annotations, and (iii) develop evaluation metrics that penalize hierarchical inconsistency. These steps could yield models that are safer and more trustworthy in real-world, hierarchy-rich settings.
Potential downsides include the amplification of existing taxonomic biases or misclassifications if the training data encodes culturally or scientifically outdated hierarchies. Researchers should therefore audit hierarchies for regional or disciplinary bias, publish data-curation protocols, and, where feasible, provide mechanisms for community feedback and correction.
Overall, we believe that exposing and remedying hierarchical blind spots in VLLMs will enable more reliable AI systems and support scientific, environmental, and industrial domains that depend on structured semantic recognition.